# Is Repeated Bayesian Interim Analysis Consequence-Free?

### v1

@@ 2.2 Decision rule and inferential metrics
Let $N$ denote the maximum sample size of the trial and $n$ an interim sample size. We consider the following Bayesian superiority stopping rule:
@@
Stop the trial and reject $H_{0}$ if $\mbox{Pr}(H_{a}|{\cal D}_{n})>C$, where $C\in(0,1)$ is a probability cutoff.
This decision rule is evaluated at each interim analysis, as well as at the end of the trial. Of note, throughout this article, we use $H_{a}$ as a plug-in term to literally represent $\theta-\theta_{0}>\delta$ (i.e., $\mbox{Pr}(H_{a}|{\cal D}_{n})\equiv\mbox{Pr}(\theta-\theta_{0}>\delta|{\cal D}_{n})$). For ease of exposition, we refer to the trial design with a fixed sample size $N$ and no interim analyses as the fixed design, and the design that incorporates interim analyses and allows early stopping as the adaptive design.
@@
To isolate the effect of interim analyses, we assume that the cutoff $C$ used in the decision rule is pre-determined based on specific criteria, such as controlling FDR or type I error rate under the fixed design. This same cutoff is applied when evaluating the operating characteristics of both the fixed and adaptive designs, ensuring that any observed differences can be attributed solely to the impact of interim analyses.
@@
To evaluate the impact of Bayesian interim analysis, we adhere to the fundamental assumption of the Bayesian framework that $\theta$ is random. This avoids the philosophical argument that evaluating Bayesian designs using the frequentist criteria (i.e., assuming $\theta$ is fixed, such as evaluating type I error at a fixed null value) is neither meaningful nor interpretable. This argument has been a main barrier to reaching a consensus on whether Bayesian methods can accommodate multiple interim looks without any consequences.
@@
Treating $\theta$ as a random variable necessitates distinguishing between the data-generating prior $\pi_{g}(\theta)$ and the user-specified prior $\pi(\theta)$ used for Bayesian inference. We assume that the data are generated according to the Bayesian process: sampling $\theta$ from $\pi_{g}(\theta)$, and generating data ${\cal D}_{n}$ based on $f(X_{i}|\theta)$. $\pi_{g}(\theta)$ can be interpreted as the distribution of the treatment effect $\theta$ across the population of all trials within a disease area, including positive and negative trials, which is generally unknown. In practice, for a specific trial, users have to specify a prior to make posterior inferences and interim decisions. They assume a user-specified prior $\pi(\theta)$ (e.g., a noninformative prior) and, conditional on the observed data ${\cal D}_{n}$, calculate $\mbox{Pr}(\theta|{\cal D}_{n})$ to make go/no-go decisions.
@@
We explore scenarios in which $\pi(\theta)$ and $\pi_{g}(\theta)$ are either identical or different; in this paper, we refer to these cases as $\pi(\theta)$ being correctly specified or misspecified, respectively. We assume that the data model $f(X_{i}|\theta)$ is always correctly specified to focus on the impact of interim analyses. To evaluate the impact of interim analyses on trial operating characteristics, we consider the following inferential quantities:
@@
  • Coverage of 95% posterior credible intervals: The probability that a 95% posterior credible interval of $\theta$ contains the true value of $\theta$, i.e., $\Pr_{\theta\sim\pi_{g},\mathcal{D}\sim f(\cdot\mid\theta)}(\theta\in CI)$, where $CI$ represents a 95% credible interval for $\theta$ derived from the posterior distribution $\pi(\theta\mid\mathcal{D})$ under $\pi(\theta)$. We consider both one-sided and symmetric credible intervals, defined as $(\theta_{0.05},\infty)$ and $(\theta_{0.025},\theta_{0.975})$, respectively, where $\theta_{\alpha}$ represents the $\alpha$-th percentile of the posterior distribution $\pi(\theta\mid\mathcal{D})$. The one-sided posterior credible interval is considered here to match the hypotheses of interest, $H_{0}:\theta-\theta_{0}<\delta$ vs $H_{a}:\theta-\theta_{0}>\delta$.
@@
  • Type I error rate A: The probability of rejecting $H_{0}$ when $H_{0}$ is true, where rejection of $H_{0}$ occurs if $\Pr(H_{a}|\mathcal{D})>C$.
  • Type I error rate B: The probability of rejecting $H_{0}$ when $H_{0}$ is true, where rejection of $H_{0}$ occurs if the one-sided posterior credible interval for $\theta$ excludes the null region (i.e., $(-\infty,\theta_{0}+\delta]$). That is, $H_{0}$ is rejected when there is no overlap between the credible interval and the null region defined by $H_{0}$.
  • Power: The probability of rejecting $H_{0}$ when $H_{0}$ is false, i.e., $\Pr(\text{Reject }H_{0}\mid\theta-\theta_{0}>\delta)$.
  • Mean sample size of the trial: The average number of patients enrolled until stopping (early or at maximum sample size).
A few remarks are warranted. First, in the definition and calculation of the above inferential quantities, $\theta$ is random. For example, when calculating Type I error rate A, $\theta$ is a random variable under $H_{0}$ (i.e., following $\pi_{g}(\theta)$ constrained by $\theta-\theta_{0}<\delta$), rather than being fixed at a point null as in the classical frequentist definition of Type I errors. This setup aims to preempt the argument that the frequentist Type I error is not meaningful within the Bayesian framework. Second, rather than relying solely on the Type I error rate, we also consider the pFDR and FDR, as one criticism of the Type I error rate is that it is conditional on the truth, which is unknown. What matters more in practice is, in positive decisions (e.g., approving a drug), how many decisions are truly positive (e.g., the approved drug is actually effective). From this perspective, the pFDR can be interpreted as the risk of regulatory regret. For this reason, in our simulation study, we choose to calibrate the cutoff $C$ using the pFDR instead of the Type I error rate. However, the specific criterion used to calibrate $C$ is not critical for our purposes, as the focus here is on evaluating the impact of incorporating interim stopping, relative to the fixed design without interim analyses. Lastly, while we use the generic notation $\theta$, in the context of RCTs with time-to-event outcomes, the bias, MSE, and coverage metrics are computed with respect to the HR, $\exp({\beta})$.
@@ 2.3 Impact of interim stopping
In this section, we examine the impact of incorporating interim analyses and early stopping on various inferential quantities, using the corresponding fixed design (with no interims) as the reference. Bayesian inference for the fixed design is well established, and using it as the reference avoids unnecessary debate over the frequentist versus Bayesian framework. This allows us to focus on whether incorporating interim analyses and stopping rules alters the operating characteristics of Bayesian inference—analogous to how interim analyses introduce multiplicity issues in the frequentist framework. Table 1 summarizes our findings, which are general and not specific to any particular model or the Bayesian decision rule discussed earlier. For example, the Bayesian decision rule previously discussed uses the same cutoff $C$ at both interim and final analyses. However, the results presented in this section also apply to scenarios where different cutoffs are used at the interim and final analyses, provided that the same final analysis cutoff is used for both the fixed and adaptive designs. In this way, the only distinction between the two designs lies in the inclusion of interim analyses.
Table 1: Impact of interim superiority stopping on inferential quantities, with respect to the fixed design without interims, assuming the same decision rule for rejecting $H_{0}$.
Quantity Interim Stopping ($\pi=\pi_{g}$) Justification Interim Stopping ($\pi\neq\pi_{g}$) Justification
Posterior mean bias Unbiased Follows from the law of total expectation under the correct prior. Biased Prior misspecification combined with selective stopping distorts the posterior mean.
FDR/pFDR Increased Multiple interim looks increase the chance of stopping on spurious signals under $H_{0}$. Often further increased Prior misspecification distorts posterior probabilities, often worsening FDR.
Type I error rate A Increased More looks at the data raise the probability of false positives. Often further increased Same mechanism, compounded by posterior miscalibration.
Type I error rate B Distorted (inflated or deflated) As its definition is conditional on $H_{0}$, $\pi_{g}$ is effectively restricted to the parameter space under $H_{0}$, leading to mismach with $\pi$. Distorted (inflated or deflated) Same mechanism, compounded by posterior miscalibration.
Not necessarily increased, as it depends on whether $\pi$ is more optimistic or pessimistic relative to $\pi_{g}$ and $H_{0}$.
@@
Let $\mathcal{D}_{n}$ be the observed data at an interim stopping time $n<N$, governed by a pre-specified stopping rule (e.g., stop when $\mbox{Pr}(H_{a}|{\cal D}_{n})>C$), and let $\hat{\theta}_{n}=\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]$ denote the posterior mean estimates based on user-specified prior $\pi(\theta)$. Under the data-generating model $\pi_{g}(\theta)$ and $f(\cdot\mid\theta)$, the bias of $\hat{\theta}_{n}$ is given by
@@
When $\pi=\pi_{g}$, by the law of total expectation (Casella and Berger, 2002), we have
@@
This means that the posterior mean under $\pi_{g}$ is an unbiased estimator of the true treatment effect, even when $\mathcal{D}_{n}$ is selected adaptively through interim analyses. In other words, interim stopping does not introduce bias when the (user-specified) prior is correctly specified (i.e., $\pi=\pi_{g}$), that is, the posterior is computed using the true prior. In Bayesian terms, this is consistent with the likelihood principle: only the observed data $\mathcal{D}_{n}$ matters for inference, not the stopping rule.
However, when $\pi\neq\pi_{g}$, $\mathbb{E}_{\pi_{g}}\left[\mathbb{E}_{\mathcal{D}_{n}\mid\theta}\left[\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\right]\right]$ generally does not equal $\mathbb{E}_{\pi_{g}}[\theta]$ because the outer expectation is under $\pi_{g}$, while the inner posterior is computed using $\pi$, and $\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\neq\mathbb{E}_{\pi_{g}}[\theta\mid\mathcal{D}_{n}]$. Interim stopping selects favorable $\mathcal{D}_{n}$ with higher-than-average treatment effects, which often amplifies prior misspecification.
@@
implying that Bayesian credible intervals maintain their nominal frequentist coverage, even under interim analyses with stopping rules (Dawid, 1982). This result follows from the law of total expectations:
@@
In contrast, when $\pi\neq\pi_{g}$, the posterior (and the resulting Bayesian credible interval) is misaligned with the data-generating process. As a result, the credible interval may under- or over-cover the true parameter, leading to deviations from nominal coverage. This is because when $\pi\neq\pi_{g}$, the posterior used to construct credible interval is misspecified, and thus the inner expectation is:
@@
FDR
Recall that the false discovery rate (FDR) is defined as
$\text{FDR}=\mathbb{E}\left[\frac{V}{R\vee 1}\right].$
Under a fixed rejection threshold $C$, conducting multiple interim analyses increases the likelihood of rejecting $H_{0}$ due to random data fluctuations—particularly early in the trial. This leads to an accumulation of false rejections across interim looks, widely known as multiplicity. In contrast, the number of true rejections under the alternative hypothesis $H_{a}$ tends to grow more slowly or even plateau. This is because the statistical power is typically high by design (e.g., 0.8), and additional interim looks offer diminishing returns in detecting true effects. As a result, the expected number of false discoveries ($V$) tends to increase more rapidly than the total number of rejections ($R$), leading to inflation of the FDR (see Supplementary Materials for a proof). This inflation occurs even when the prior distribution is correctly specified (i.e., $\pi=\pi_{g}$), and may be further exacerbated under prior misspecification (i.e., $\pi\neq\pi_{g}$). A similar argument applies to the pFDR as well.
Type I error rates A and B
Bayesian interim analyses inflate type I error rate A, regardless of whether $\pi(\theta)$ is correctly or incorrectly specified. This is because at each interim analysis, there is a non-zero probability that the posterior probability exceeds the threshold, i.e., $\Pr(H_{a}\mid\mathcal{D}_{n})>C$, purely due to random fluctuations in the observed data. That is, $H_{0}$ may be falsely rejected by chance when $\theta-\theta_{0}<\delta$, resulting in inflation of the type I error rate A. This inflation is often further exacerbated under prior misspecification (i.e., $\pi\neq\pi_{g}$), as stopping rules tend to preferentially select data with extreme or favorable outcomes (e.g., indicating early efficacy), thereby biasing the posterior toward an overly optimistic direction.
Since Type I error rate B is defined based on credible intervals, one might expect that interim stopping would not inflate it when $\pi=\pi_{g}$. However, this is not the case. Because Type I error rate B is defined conditional on $H_{0}$, not all parameter values generated under $\pi_{g}$ contribute to its calculation. In effect, only the portion of $\pi_{g}$ that lies within the null parameter space (e.g., $\theta-\theta_{0}<\delta$) is relevant. This creates a mismatch between the data-generating prior $\pi_{g}$ and the prior $\pi$ used for inference, even when they are nominally the same. As a result, Type I error rate B can differ from that of a fixed design, despite using $\pi=\pi_{g}$. This deviation becomes more pronounced when $\pi\neq\pi_{g}$.

### v2

@@ 2.2 Decision rule and inferential metrics
Let $N$ denote the maximum sample size of the trial and $n$ an interim sample size. Let $\theta_{0}$ denote the treatment effect for the control. In single-arm trials, $\theta_{0}$ is assumed to be a known constant, whereas in two-arm RCTs, $\theta_{0}$ is treated as an unknown parameter. The treatment is considered effective, relative to the control, if $\theta>\theta_{0}$.
We consider the following Bayesian superiority decision rule, evaluated at each interim analysis and, if the trial is not stopped early, again at the final analysis:
@@
Stop the trial and claim that the treatment is effective if $\mbox{Pr}(\theta>\theta_{0}\,|\,{\cal D}_{n})>C$, where $C\in(0,1)$ is a probability cutoff.
For ease of exposition, we refer to the trial design with a fixed sample size $N$ and no interim analyses as the fixed design, and the design that incorporates interim analyses and allows early stopping as the adaptive design. For notational simplicity, we assume hereafter that $\theta_{0}$ is a constant. However, all assessments and results presented later also apply to settings in which $\theta_{0}$ is random, as in RCTs.
@@
To isolate the effect of interim analyses, we apply the same decision rule when evaluating the operating characteristics of both the fixed and adaptive designs. In other words, we do not calibrate the cutoff $C$ separately for the two designs. This ensures that any observed differences can be attributed solely to the presence of interim analyses.
@@
In addition, in our assessment, we adhere to the fundamental assumption of the Bayesian framework that $\theta$ is random. This avoids the philosophical argument that evaluating Bayesian designs using the frequentist criteria (i.e., assuming $\theta$ is fixed, such as evaluating type I error at a fixed null value) is neither meaningful nor interpretable. This argument has been a main barrier to reaching a consensus on whether Bayesian methods can accommodate multiple interim looks without any consequences.
@@
Treating $\theta$ as a random variable necessitates distinguishing between the data-generating prior $\pi_{g}(\theta)$ and the user-specified prior $\pi(\theta)$ used for Bayesian inference. We assume that the data are generated according to the Bayesian process: sampling $\theta$ from $\pi_{g}(\theta)$, and generating data ${\cal D}_{n}$ based on $f(X_{i}|\theta)$. $\pi_{g}(\theta)$ can be interpreted as the distribution of the treatment effect $\theta$ across the population of all trials within a therapeutic area, including positive and negative trials, which is generally unknown. In practice, $\pi_{g}(\theta)$ is typically unknown. For a specific trial, users have to specify a prior to make posterior inferences and interim decisions. They assume a user-specified prior $\pi(\theta)$ (e.g., a noninformative prior) and, conditional on the observed data ${\cal D}_{n}$, calculate $\mbox{Pr}(\theta|{\cal D}_{n})$ to make decisions.
@@
We explore scenarios in which $\pi(\theta)$ and $\pi_{g}(\theta)$ are either identical or different; in this paper, we refer to these cases as $\pi(\theta)$ being correctly specified or misspecified, respectively. We assume that the data model $f(X_{i}|\theta)$ is always correctly specified to focus on the impact of interim analyses. To evaluate the impact of interim analyses on trial operating characteristics, we consider the following key performance metrics:
@@
  • Coverage of 95% posterior credible intervals: The probability that a 95% posterior credible interval of $\theta$ contains the true value of $\theta$, i.e., $\Pr_{\theta\sim\pi_{g},\mathcal{D}\sim f(\cdot\mid\theta)}(\theta\in CI)$, where $CI$ represents a 95% credible interval for $\theta$ derived from the posterior distribution $\pi(\theta\mid\mathcal{D})$ under $\pi(\theta)$. Specifically, in our simulation, we consider symmetric credible intervals $(\theta_{0.025},\theta_{0.975})$, where $\theta_{\alpha}$ represents the $\alpha$-th percentile of the posterior distribution $\pi(\theta\mid\mathcal{D})$.
@@
    Although the FDR was originally developed as a tool to control frequentist error rates, \@BBOPcitet\@BAP\@BBNStorey03\@BBCP noted that it is inherently more Bayesian in nature, since it is defined conditional on the observed data, in contrast to the type I error rate, which is defined conditional on the true parameter $\theta$.
    The FDR/pFDR is of substantial practical interest. If we assume that positive trials lead to drug approvals, then the pFDR quantifies, among all approval decisions, the proportion that are incorrect, that is, cases in which the approved drug is actually not effective. From this perspective, the pFDR can be interpreted as the risk of regulatory regret.
  • ATIE (Average Type I Error Rate): The probability of claiming the treatment effective when it is actually not effective (i.e., $\theta\leq\theta_{0}$). Formally, ATIE $\mathbb{E}_{\theta\sim\pi_{g},\mathcal{D}\sim f(\cdot\mid\theta)}[{\cal S}\,|\,\theta\leq\theta_{0}]$, where ${\cal S}=I\{\mbox{Pr}(\theta>\theta_{0}\,|\,{\cal D}_{n})>C\}$. The ATIE can be regarded as the Bayesian counterpart of the frequentist type I error rate, as it averages over the distribution of the $\theta$ in the region $\theta\leq\theta_{0}$ (i.e., when the treatment is not effective), treating $\theta$ as a random variable rather than a fixed parameter. The ATIE is sometimes referred to as the Bayesian type I error rate \@BBOPcitep\@BAP\@BBN(best2025beyond)\@BBCP, and is described in the FDA’s Guidance for the Use of Bayesian Statistics in Medical Device Clinical Trials as an alternative definition of type I error for Bayesian methods \@BBOPcitep\@BAP\@BBN(FDA10)\@BBCP.
A few remarks are warranted. First, in the definition and calculation of the above performance metrics, $\theta$ is random. This setup aims to preempt the argument that the evaluation is not consistent with the Bayesian framework. Second, our focus here is solely on evaluating the impact of repeated interim analyses on these commonly used performance metrics. It is not our goal to argue for or against the importance or appropriateness of any particular metric, especially from a philosophical standpoint (e.g., some may argue that ATIE conditions on $\theta$, whereas for a given trial Bayesians are primarily interested in conditioning on the observed data $\cal D$}). Rather, users can decide for themselves how relevant a given metric is to their specific application. Lastly, while we use the generic notation $\theta$, in the context of RCTs with time-to-event outcomes, the inferential quantities are computed with respect to the HR, $\exp({\beta})$.
@@ 2.3 Impact of interim stopping
In this section, we examine the impact of incorporating interim analyses and early superiority stopping on various design performance metrics, using the corresponding Bayesian fixed design (with no interims) as the reference. In this way, the only distinction between the two designs lies in the inclusion of interim analyses. This allows us to focus on whether incorporating interim analyses and stopping rules alters the operating characteristics of Bayesian inference, avoiding unnecessary debate over the frequentist versus Bayesian framework. Table 1 summarizes our findings, which are general and not specific to any particular model or the Bayesian decision rule discussed earlier.
@@
Let $\mathcal{D}_{n}$ be the observed data at an interim stopping time $n<N$, governed by a pre-specified stopping rule (e.g., stop when $\mbox{Pr}(\theta>\theta_{0}\,|\,{\cal D}_{n})>C$), and let $\hat{\theta}_{n}=\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]$ denote the posterior mean estimates based on user-specified prior $\pi(\theta)$. Under the data-generating model $\pi_{g}(\theta)$ and $f(\cdot\mid\theta)$, the bias of $\hat{\theta}_{n}$ is given by
@@
When $\pi=\pi_{g}$, by the law of total expectation \@BBOPcitep\@BAP\@BBN(Casella02)\@BBCP, we have
@@
This means that the posterior mean under $\pi_{g}$ is an unbiased estimator of the true treatment effect, even when $\mathcal{D}_{n}$ is selected adaptively through interim analyses. In other words, interim stopping does not introduce bias when the (user-specified) prior is correctly specified (i.e., $\pi=\pi_{g}$). In Bayesian terms, this is consistent with the likelihood principle: only the observed data $\mathcal{D}_{n}$ matters for inference, not the stopping rule.
This result has been cited in some of the literature to support the claim that Bayesian analysis and stopping can be performed at any point during a trial without adjustment. However, what is often not explicitly highlighted is that this property holds only when the user-specified prior used for inference and decision-making is identical to the underlying data-generating prior. One goal of this article is to clarify this distinction.
Unfortunately, when $\pi\neq\pi_{g}$, which is the dominant case in practice, $\mathbb{E}_{\pi_{g}}\left[\mathbb{E}_{\mathcal{D}_{n}\mid\theta}\left[\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\right]\right]$ generally does not equal $\mathbb{E}_{\pi_{g}}[\theta]$ because the outer expectation is under $\pi_{g}$, while the inner posterior is computed using $\pi$, and $\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\neq\mathbb{E}_{\pi_{g}}[\theta\mid\mathcal{D}_{n}]$. Interim stopping selects favorable $\mathcal{D}_{n}$ with higher-than-average treatment effects, which often amplifies prior misspecification, resulting in biased posterior mean estimate.
@@
implying that Bayesian credible intervals maintain their nominal frequentist coverage, even under interim analyses with stopping rules \@BBOPcitep\@BAP\@BBN(Dawid82)\@BBCP. This result follows from the law of total expectations:
@@
However, when $\pi\neq\pi_{g}$, the posterior (and the resulting Bayesian credible interval) is misaligned with the data-generating process. As a result, the credible interval may under- or over-cover the true parameter, leading to deviations from nominal coverage. This is because when $\pi\neq\pi_{g}$, the posterior used to construct credible interval is misspecified, and thus the inner expectation is:
@@
This result further demonstrates that the claim that Bayesian analysis and stopping can be performed at any point during a trial without adjustment is neither entirely true nor entirely false. It holds only when the user-specified prior used for inference and decision-making is identical to the underlying data-generating prior, a condition that rarely, if ever, holds in practice. This finding reconciles the conflicting arguments regarding Bayesian multiplicity that have long persisted in the literature.
FDR/pFDR
Bayesian interim analyses and stopping generally lead to inflation of the FDR and pFDR for the following reason. Recall that the FDR is defined as $\text{FDR}=\mathbb{E}\left[\frac{V}{R\vee 1}\right].$ Under a fixed rejection threshold $C$, conducting multiple interim analyses increases the likelihood of $\mbox{Pr}(\theta>\theta_{0}\,|\,{\cal D}_{n})>C$ due to random data fluctuations, particularly early in the trial. This leads to an accumulation of false discovery across interim looks, widely known as multiplicity. In contrast, the number of true discovery when the treatment is truly effective tends to grow more slowly or even plateau. This is because the statistical power is typically high by design (e.g., 0.8), and additional interim looks offer diminishing returns in detecting true effects. As a result, the expected number of false discoveries ($V$) tends to increase more rapidly than the total number of discoveries ($R$), leading to inflation of the FDR (see Supplementary Materials for a proof). This inflation occurs even when the prior distribution is correctly specified (i.e., $\pi=\pi_{g}$), and may be further exacerbated under prior misspecification (i.e., $\pi\neq\pi_{g}$). A similar argument applies to the pFDR as well.
ATIE
Bayesian interim analyses inflate ATIE, regardless of whether $\pi(\theta)$ is correctly or incorrectly specified. This is because at each interim analysis, there is a non-zero probability that the posterior probability exceeds the threshold, i.e., $\Pr(\theta>\theta_{0}\mid\mathcal{D}_{n})>C$, purely due to random fluctuations in the observed data, when $\theta\leq\theta_{0}$. This resulting in inflation of ATIE. This inflation is often further exacerbated under prior misspecification (i.e., $\pi\neq\pi_{g}$), as stopping rules tend to preferentially select data with extreme or favorable outcomes (e.g., indicating early efficacy), thereby biasing the posterior toward an overly optimistic direction.